Alice is ten years less than twice as old as Bonnie. If the sum of their ages is 26 years, find their ages.
Answer by CharlesG2(834)

Alice is ten years less than twice as old as Bonnie. If the sum of their ages is 26 years, find their ages.
A Alice, B Bonnie
A=2B-10
A+B=26
A=26-B
26-B=2B-10
26=3B-10
36=3B
12=B
A+12=26
A=14
Alice is 14 and Bonnie is 12